The Minute Stop Market has 1 pump for gasoline, which can service 10 customers per hour (Poisson distributed). Cars arrive at the pump at a rate of 5 per hour (Poisson distributed). Determine the average queue length, the average time a car is in the system, and the average time a car must wait.
Answer #1

Service Rate(μ) = 10 cares /hr

λ = 5 cars/hr

1 Average queue length = μ/(μ-λ)

= 10/(10-5)

= 2

2 Average time a car is in the system = 1/(μ-λ)

= 1/(10-5)

= 0.2 hrs

= 0.2*60

= 12min

3 The average time a car must wait = λ/μ(μ-λ)

= 5/10(10-5)

= 0.1 hrs

= 0.1*60= 6min